# Rethinking Audio-Visual Adversarial Vulnerability from Temporal and Modality Perspectives

### v2

@@ 1 Introduction
Audio-visual models, capable of integrating both auditory and visual information, have gained significant traction in recent years due to their ability to create a comprehensive understanding of the surrounding world (Zhu et al., 2021a; Wei et al., 2022; Li et al., 2022a). These models have demonstrated remarkable success in a wide range of applications, including multimedia analysis (Dimoulas, 2016), human-computer interaction (Zhen et al., 2023), and autonomous systems (Guo et al., 2023). However, a critical challenge lies in their susceptibility to adversarial attacks. These attacks can craft imperceptible perturbations to the input data, causing audio-visual models to make erroneous predictions (Wang et al., 2022) or interpretations (Han et al., 2023). Such errors can have disastrous consequences, especially in safety-critical domains like auto-driving (Kloukiniotis et al., 2022) and identity verification (Zhang et al., 2021).
@@ 2.1 Audio-visual Learning
The field of audio-visual learning encompasses a wide range of tasks, including audio-visual event recognition (Brousmiche et al., 2019; Xia & Zhao, 2022; Brousmiche et al., 2022), separation (Wu et al., 2019a; Majumder et al., 2021; Majumder & Grauman, 2022), localization (Wu et al., 2019b, 2021; Mo & Morgado, 2022), correspondence learning (Min et al., 2020; Zhu et al., 2021b; Morgado et al., 2021), representation learning (Zhou et al., 2019; Cheng et al., 2020; Rahman et al., 2021), and cross-modal generation (Chen et al., 2017; Hao et al., 2018; Sung-Bin et al., 2023). Among these, audio-visual event recognition stands out as a fundamental task (Gao et al., 2024) that has attracted significant research attention, particularly regarding robustness and security issues (Yang et al., 2023b).
@@ 4.2 Modality Misalignment-based Attack
Table 1: The average black-box attack success rate (A.S.R.) and cosine similarity (C.S.) of the audio-visual adversarial examples generated by different attack methods.
Method FGSM I-FGSM MI-FGSM NI-FGSM
A.S.R. (%) $31.7$ $30.2$ $53.1$ $54.8$
C.S. (%) $47.6$ $49.3$ $35.2$ $34.9$
From the results shown in table 1, we can see that as the attack success rate against black-box models (i.e., adversarial transferability) increases, the cosine similarity between audio and visual feature vectors decreases. This correlation reinforces our claim that disrupting semantic alignment between modalities enhances transferability.
@@ 5.3 Our Approach
Adversarial curriculum training. Previous studies (Yu et al., 2022; Kim et al., 2021; Rice et al., 2020) have identified that adversarial training easily gets overfitted to certain attack methods and settings, leaving the model vulnerable to others. To address this issue and improve the generalization ability, we propose a randomized adversarial curriculum learning to optimize eq. 5.
Concretely, our proposed randomized adversarial curriculum learning approach incorporates two strategies:
@@
The curriculum learning approach hinges on a cyclical variation in the data-level masking ratio $\rho_{x}$ and the model-level dropout ratio $\rho_{f}$. This cyclic design offers a key advantage: it gradually increases the training difficulty. During the initial stages of the cycle, both sampling and dropout ratios are lower and the number of steps to generate attacks is reduced. This translates to a simpler training task with fewer adversarial perturbations applied to the data. As both ratios and the number of steps gradually increase, complex adversarial examples are created, resulting in a more difficult learning task. With this cyclic design, the models are trained on diverse perturbations, thus fully exploring the loss landscape for the searching minimum and improving the robustness against different attacks.
@@ 6.1 Experimental Setup
Datasets. We use the Kinetics-Sounds (Arandjelovic & Zisserman, 2017) for evaluation, which contains $1,551,610$ second video clips in $27$ human action categories. We split the dataset into $7:2:1$ for training, validation, and testing. We also conduct experiments on MIT-MUSIC (Zhao et al., 2018b) for further verification, which is provided in the appendix.
@@ 6.5 Ablation Study
On the sampling ratio for the adversarial training. In our approaches, we propose to sample a ratio of frames to generate the temporal-universal adversarial perturbation for efficient adversarial training. There is a balance between the time consumption of adversarial training and the adversarial robustness by selecting different sample ratios. Here, we conduct an ablation study on the sampling ratio. As shown in table 2, with increasing the sampling ratio in adversarial training, the defense performance is improved, but it also leads to more time consumption, supporting our argument on the balance. We advocate the selection of $15\%$ in adversarial training. At the same time, a larger sampling ratio does not improve the adversarial robustness much but introduces more time consumption, and a smaller sampling ratio harms defense performance.
Table 2: Ablation study of the sampling ratio on the performance of adversarial training. We use our proposed TMA to generate the adversarial example for adversarial training.
Sampling ratio 5% 10% 15% 20% 25% 30% 35% 40%
Successful defense rate (%) 72.1 75.4 81.5 82.0 82.6 82.9 83.0 83.0
Training time (Hours) 6.2 9.3 17.8 18.9 21.3 26.4 31.8 34.2

### v3

@@ 1 Introduction
Audio-visual models, capable of integrating both auditory and visual information, have gained significant traction in recent years due to their ability to create a comprehensive understanding of the surrounding world (Li et al., 2022a; Song et al., 2023; 2024a; 2024b). These models have demonstrated remarkable success in a wide range of applications, including multimedia analysis (Dimoulas, 2016), human-computer interaction (Zhen et al., 2023), and autonomous systems (Guo et al., 2023). However, a critical challenge lies in their susceptibility to adversarial attacks (Zhang et al., 2024b; a; Yao et al., 2024). These attacks can craft imperceptible perturbations to the input data, causing audio-visual models to make erroneous predictions (Wang et al., 2022) or interpretations (Han et al., 2023). Such errors can have disastrous consequences, especially in safety-critical domains like auto-driving (Kloukiniotis et al., 2022) and identity verification (Zhang et al., 2021).
@@ 2.1 Audio-visual Learning
The field of audio-visual learning encompasses a wide range of tasks, including audio-visual event recognition (Brousmiche et al., 2019; Xia & Zhao, 2022; Brousmiche et al., 2022), separation (Wu et al., 2019a; Majumder et al., 2021; Majumder & Grauman, 2022; Huang et al., 2023a), localization (Wu et al., 2019b; 2021; Huang et al., 2023b), correspondence learning (Min et al., 2020; Zhu et al., 2021; Morgado et al., 2021), representation learning (Zhou et al., 2019; Cheng et al., 2020; Rahman et al., 2021), and cross-modal generation (Chen et al., 2017; Hao et al., 2018; Sung-Bin et al., 2023; Liang et al., 2023a; 2024; Huang et al., 2024; Liang et al., 2023b). Among these, audio-visual event recognition stands out as a fundamental task (Gao et al., 2024) that has attracted significant research attention, particularly regarding robustness and security issues (Yang et al., 2023b).
@@ 4.2 Modality Misalignment-based Attack
Figure 3: The average black-box attack success rate (A.S.R.) and cosine similarity (C.S.) of the audio-visual adversarial examples generated by different attacks.
From the results shown in fig. 3, we can see that as the attack success rate against black-box models (i.e., adversarial transferability) increases, the cosine similarity between audio and visual feature vectors decreases. This correlation reinforces our claim that disrupting semantic alignment between modalities enhances transferability.
@@ 5.3 Our Approach
Adversarial curriculum training. Previous studies (Yu et al., 2022; Kim et al., 2021; Rice et al., 2020) have identified that adversarial training easily gets overfitted to certain attack methods and settings, leaving the model vulnerable to others. To address this issue and improve the generalization ability, we propose a randomized adversarial curriculum learning to optimize eq. 5. Concretely, our proposed randomized adversarial curriculum learning approach incorporates two strategies:
@@
The curriculum learning approach hinges on a cyclical variation in the data-level masking ratio $\rho_{x}$ and the model-level dropout ratio $\rho_{f}$. This cyclic design offers an advantage: it gradually increases the training difficulty. During the initial stages of the cycle, both sampling and dropout ratios are lower and the number of steps to generate attacks is reduced. This translates to a simpler training task with fewer adversarial perturbations applied to the data. As both ratios and the number of steps gradually increase, complex adversarial examples are created, resulting in a more difficult learning task. With this cyclic design, the models are trained on diverse perturbations, thus fully exploring the loss landscape for the searching minimum and improving the robustness against attacks.
@@ 6.1 Experimental Setup
Datasets. We use the Kinetics-Sounds (Arandjelovic & Zisserman, 2017) for evaluation, which contains $1,551,610$ second video clips in $27$ human action categories. For model training, we split the dataset into $7:2:1$ for training, validation, and testing. We also conduct experiments on MIT-MUSIC (Zhao et al., 2018b) for further verification, which is provided in the appendix.
@@ 6.5 Ablation Study
Figure 9: Ablation study of the sampling ratio on the performance of adversarial training. We use our proposed TMA to generate the adversarial example for adversarial training.
On the sampling ratio for the adversarial training. In our approaches, we propose to sample a ratio of frames to generate the temporal-universal adversarial perturbation for efficient adversarial training. There is a balance between the time consumption of adversarial training and the adversarial robustness by selecting different sample ratios. Here, we conduct an ablation study on the sampling ratio. As shown in fig. 9, with increasing the sampling ratio in adversarial training, the defense performance is improved, but it also leads to more time consumption, supporting our argument on the balance. At the same time, a larger sampling ratio does not improve the adversarial robustness much but introduces more time consumption, and a smaller sampling ratio harms defense performance.